# Embracing Dynamics: Dynamics-aware 4D Gaussian Splatting SLAM

### v1

@@ Abstract
Simultaneous localization and mapping (SLAM) technology now has photorealistic mapping capabilities thanks to the real-time high-fidelity rendering capability of 3D Gaussian splatting (3DGS). However, due to the static representation of scenes, current 3DGS-based SLAM encounters issues with pose drift and failure to reconstruct accurate maps in dynamic environments. To address this problem, we present D4DGS-SLAM, the first SLAM method based on 4DGS map representation for dynamic environments. By incorporating the temporal dimension into scene representation, D4DGS-SLAM enables high-quality reconstruction of dynamic scenes. Utilizing the dynamics-aware InfoModule, we can obtain the dynamics, visibility, and reliability of scene points, and filter stable static points for tracking accordingly. When optimizing Gaussian points, we apply different isotropic regularization terms to Gaussians with varying dynamic characteristics. Experimental results on real-world dynamic scene datasets demonstrate that our method outperforms state-of-the-art approaches in both camera pose tracking and map quality.
@@ I INTRODUCTION
Simultaneous Localization and Mapping (SLAM) has been a cornerstone in robotics and computer vision, enabling applications such as autonomous navigation, augmented reality, and 3D reconstruction. SLAM is a crucial technology for robotics and computer vision. Recently, SLAM has benefited from 3D Gaussian Splatting (3DGS) [7], which offers explicit map representation and real-time high-fidelity rendering. 3DGS-based methods such as SplaTAM [6], MonoGS [9], and GS-SLAM [24] have achieved state-of-the-art (SOTA) map reconstruction quality and speed in static environments. However, these methods face challenges in dynamic environments with moving objects, leading to camera pose estimation drift and the appearance of artifacts in the map, or even map reconstruction failure, as shown in Fig 1. Although recent work like DG-SLAM [24] attempts to address tracking issues in dynamic environments by filtering dynamic objects, it still relies on a static 3DGS representation, which cannot fully reconstruct the complete map.
@@
To achieve accurate camera pose estimation and real-time high-fidelity map reconstruction in dynamic scenes, we propose D4DGS-SLAM. To the best of our knowledge, this is the first SLAM system based on dynamic map representation. Unlike existing 3DGS-based SLAM methods that filter out dynamic objects, we choose to “embrace” dynamics by using 4D Gaussian Splatting (4DGS) [26] with a temporal dimension for map representation and rendering. 4DGS enables our system to effectively learn the spatio-temporal characteristics of the scene, thereby reconstructing a map with few artifacts without filtering out dynamic objects. To reduce tracking drifts in dynamic environments when using photometric and geometric error-based optimization, we introduce a long-term effective any point tracking (LEAP) [3] for dynamics-aware tracking point filtering. By leveraging the dynamics and reliability of anchors in the scene provided by LEAP, we can select stable and static points in the current frame for tracking, significantly improving the accuracy of camera pose estimation. Dynamic awareness also facilitates the densification and pruning of Gaussians over time, allowing the Gaussian set to better fit dynamic scenes. The experimental results on real-world dynamic datasets demonstrate that our method achieves SOTA performance in both camera pose estimation and mapping quality. Our contributions are summarized as follows:
@@
    We enhance the SLAM system with dynamic awareness, providing strategies and a pipeline for tightly integrating dynamic information, which significantly improves its tracking and mapping capabilities.
@@ II Related Work
3D Gaussian Splatting (3DGS) [7] revolutionized scene representation through explicit, differentiable rasterization. By modeling scenes as anisotropic Gaussians with view-dependent appearance, 3DGS achieves photorealistic rendering at 200+ FPS while maintaining explicit geometric control. This breakthrough inspired Gaussian-based SLAM systems like SplaTAM [6], MonoGS [9] and GS-SLAM [24], which optimize Gaussian parameters alongside camera poses. Though these methods set new SOTA in static scene reconstruction, they inherit the static assumption bottleneck, namly dynamic objects cause irreversible corruption of the Gaussian map due to unmodeled temporal variations. Despite efforts by recent works such as DG-SLAM [24] to tackle tracking issues in dynamic environments by filtering out dynamic objects, the reliance on a static 3DGS representation limits its ability to fully reconstruct an accurate map. Unlike the aforementioned works, our approach embraces dynamic objects and learns their characteristics rather than removing them. As demonstrated below, this concept significantly enhances the capabilities of the SLAM system in dynamic environments.
@@ III-B Long-term Effective Any Point Tracking
Considering a sequence of $S$ consecutive RGB images $\mathbf{I}=[\mathbf{I}_{1},...,\mathbf{I}_{S}]$, LEAP predicts the trajectory $\mathbf{X}=[\mathbf{x}_{1},...,\mathbf{x}_{S}]$, $\mathbf{M_{d}}=[\mathbf{m_{d}}_{1},...,\mathbf{m_{d}}_{S}]$, visibility $\mathbf{V}=[v_{1},...,v_{S}]$ and reliability of visibility $\mathbf{\Phi}=[\phi_{1},...,\phi_{S}]$ for a query point $\mathbf{x}_{q}$ in frame $s_{q}$. This can is expressed as:
@@
Figure 4: Visualization of anchors with high dynamics (in red) and low reliability (in yellow)
Fig. 4 illustrates the highly dynamic points and points with uncertain dynamic status in the current frame, as identified by the LEAP module. This visualization highlights the ability of LEAP to distinguish between dynamic and uncertain regions, which is crucial for accurate tracking and mapping in dynamic environments. Unlike other pairwise trackers, LEAP is capable of long-term point tracking across image sequences, demonstrating SOTA performan in dynamic environments. This capability ensures its effectiveness in camera pose estimation and dynamic environment mapping within SLAM systems. Our experimental results further validate this claim.
@@ V CONCLUSIONS
In this paper, we present D4DGS-SLAM, the first SLAM method based on dynamic map representation and dynamics-aware module. Experimental results on dynamic datasets demonstrate that our method effectively captures the dynamic characteristics of the scene, outperforming state-of-the-art algorithms in both camera tracking accuracy and mapping quality. In addition to introducing 4DGS into map representation for the first time, we also provide a framework that integrates dynamic information into 4D Gaussian optimization. This will allow our algorithm to improve as dynamic point tracking methods advance. In the future, our work will focus on exploring how to use fewer Gaussians to represent scenes, enabling the deployment of our algorithm on small robots such as drones.

### v2

@@ Abstract
Simultaneous localization and mapping (SLAM) technology has recently achieved photorealistic mapping capabilities thanks to the real-time, high-fidelity rendering enabled by 3D Gaussian Splatting (3DGS). However, due to the static representation of scenes, current 3DGS-based SLAM encounters issues with pose drift and failure to reconstruct accurate maps in dynamic environments. To address this problem, we present D4DGS-SLAM, the first SLAM method based on 4DGS map representation for dynamic environments. By incorporating the temporal dimension into scene representation, D4DGS-SLAM enables high-quality reconstruction of dynamic scenes. Utilizing the dynamics-aware InfoModule, we can obtain the dynamics, visibility, and reliability of scene points, and filter out unstable dynamic points for tracking accordingly. When optimizing Gaussian points, we apply different isotropic regularization terms to Gaussians with varying dynamic characteristics. Experimental results on real-world dynamic scene datasets demonstrate that our method outperforms state-of-the-art approaches in both camera pose tracking and map quality.
@@ I INTRODUCTION
Simultaneous Localization and Mapping (SLAM) has been a cornerstone in robotics and computer vision, enabling applications such as autonomous navigation, augmented reality, and 3D reconstruction. Recently, SLAM has benefited from 3D Gaussian Splatting (3DGS) [7], which offers explicit map representation and real-time high-fidelity rendering. 3DGS-based methods such as SplaTAM [6], MonoGS [9], and GS-SLAM [24] have achieved state-of-the-art (SOTA) map reconstruction quality and speed in static environments. However, these methods face challenges in dynamic environments with moving objects, leading to camera pose estimation drift and the appearance of artifacts in the map, or even to map reconstruction failure, as shown in Fig. 1. Recent work such as DG-SLAM [24] improves tracking performance in dynamic environments by filtering dynamic objects. However, since it depends on a static 3DGS representation, it cannot fully reconstruct the entire scene nor produce high-quality complete maps.
@@
To achieve accurate camera pose estimation and real-time high-fidelity map reconstruction in dynamic scenes, we propose DD4DGS-SLAM. To the best of our knowledge, this work presents the first SLAM system based on a dynamic map representation. Unlike existing 3DGS-based SLAM methods that filter out dynamic objects, we choose to “embrace” dynamics by using 4D Gaussian Splatting (4DGS) [26] with a temporal dimension for map representation and rendering. 4DGS enables our system to effectively learn the spatio-temporal characteristics of the scene, thereby reconstructing a map with few artifacts without filtering out dynamic objects. To reduce tracking drifts in dynamic environments when using photometric and geometric error-based optimization, we introduce a long-term effective any point tracking (LEAP) [3] for dynamics-aware tracking point filtering. By leveraging the dynamics and reliability of anchors in the scene provided by LEAP, we can select stable and static points in the current frame for tracking, significantly improving the accuracy of camera pose estimation. Dynamic awareness also facilitates the densification and pruning of Gaussians over time, allowing the Gaussian set to better fit dynamic scenes. The experimental results on real-world dynamic datasets demonstrate that our method achieves SOTA performance in both camera pose estimation and mapping quality. Our contributions are summarized as follows:
@@
    We enhance the SLAM system with the dynamics-aware InfoModule, upon which we build strategies and a pipeline for tightly integrating dynamic information, significantly improving tracking and mapping capabilities.
@@ II Related Work
3DGS [7] revolutionized scene representation through explicit, differentiable rasterization. By modeling scenes as anisotropic Gaussians with view-dependent appearance, 3DGS achieves photorealistic rendering at 200+ FPS while maintaining explicit geometric control. This breakthrough inspired Gaussian-based SLAM systems like SplaTAM [6], MonoGS [9] and GS-SLAM [24], which optimize Gaussian parameters alongside camera poses. Though these methods set new SOTA in static scene reconstruction, they inherit the static assumption bottleneck, namly dynamic objects cause irreversible corruption of the Gaussian map due to unmodeled temporal variations. Despite efforts by recent works such as DG-SLAM [24] to tackle tracking issues in dynamic environments by filtering out dynamic objects, the reliance on a static 3DGS representation limits its ability to fully reconstruct an accurate map. Unlike the aforementioned works, our approach embraces dynamic objects and learns their characteristics rather than removing them. As demonstrated below, this concept significantly enhances the capabilities of the SLAM system in dynamic environments.
@@ III-B Long-term Effective Any Point Tracking
Considering a sequence of $S$ consecutive RGB images $\mathbf{I}=[\mathbf{I}_{1},...,\mathbf{I}_{S}]$, LEAP predicts the trajectory $\mathbf{X}=[\mathbf{x}_{1},...,\mathbf{x}_{S}]$, dynamic track label $\mathbf{M_{d}}=[\mathbf{m_{d}}_{1},...,\mathbf{m_{d}}_{S}]$, visibility $\mathbf{V}=[v_{1},...,v_{S}]$ and reliability of visibility $\mathbf{\Phi}=[\phi_{1},...,\phi_{S}]$ for a query point $\mathbf{x}_{q}$ in frame $s_{q}$. This can is expressed as:
@@
Figure 4: Anchors with high dynamics (in red) and low reliability (in yellow)
Fig. 4 illustrates the highly dynamic points and points with uncertain dynamic status in the current frame, as identified by the LEAP module. This visualization highlights the ability of LEAP to distinguish between dynamic and uncertain regions, which is crucial for accurate tracking and mapping in dynamic environments. Unlike other pairwise trackers, LEAP is capable of long-term point tracking across image sequences, demonstrating SOTA performance in dynamic environments. This capability ensures its effectiveness in camera pose estimation and dynamic environment mapping within SLAM systems. Our experimental results further validate this claim.
@@ V CONCLUSIONS
In this paper, we present D4DGS-SLAM, the first SLAM method based on dynamic map representation and dynamics-aware module. Experimental results on dynamic datasets demonstrate that our method effectively captures the dynamic characteristics of the scene, outperforming state-of-the-art algorithms in both camera tracking accuracy and mapping quality. In addition to introducing 4DGS into map representation for the first time, we also provide a framework that integrates dynamic information into 4D Gaussian optimization. This will allow our algorithm to improve as dynamic point tracking methods advance. Nonetheless, this work serves as a preliminary investigation, with many aspects warranting further study and the potential to reveal additional challenges. For example, tracking performance may degrade under heavy occlusion by dynamic objects, due to inherent difficulties in distinguishing between dynamic and static scene components. Moreover, the current reliance on a large number of Gaussians to represent scenes restricts the applicability of our algorithm on resource-constrained platforms, such as small robots or drones. Future work will aim to explore these and other limitations in greater depth, and to develop effective solutions accordingly.